# AS3D: 2D-Assisted Cross-Modal Understanding with Semantic-Spatial Scene Graphs for 3D Visual Grounding

### v2

@@ IV-E1 Analysis of Key Components
TABLE VI: Ablation studies of graph attention layers on Nr3D
@@
TABLE VII: Ablation studies of various CLIP models on Nr3D
@@
IV-E2 Losses of 2D-Assisted Training
In Section III-A1, we employ various contrastive losses to define the 3D object-text alignment loss. The losses that leverage 2D pre-trained multi-modal features play an important role in 3D vision-language learning. Table IV presents the quantitative results of the impact of various loss functions on localization performance. “$PI^{*}$” denotes the contrastive loss between the point cloud features and CLIP-extracted 2D features. “$TT^{*}$” represents the contrastive loss between text features extracted by our framework and CLIP-generated textual embeddings. “$IC$” refers to the interactive contrastive loss that aligns 3D and 2D visual features with the interchangeable corresponding text features. The tabular data demonstrates that both text contrastive and interaction contrastive losses substantially improve 3D multi-modal learning. This suggests that incorporating 2D-assisted training can help mitigate the issue of cross-modality understanding in situations where 3D data is limited.
IV-E3 Structure of Object Encoder
The object encoder based on the PointNet++ network is a layer-by-layer process of geometric sampling and feature aggregation, so at which layer is it optimal to fuse 2D pre-trained attributes and generate the new encoding branch? Table V illustrates the final localization effect on the set of attribute-guided encoding branches across different layers. “layer-$i$” denotes that the network architecture bifurcates into two parallel branches at the $i$-th layer, each dedicated to processing distinct feature representations. “att-dim‘’ denotes the mapping dimension of the 2D object attribute features extracted from CLIP before fusion with the point cloud features, to match the output dimension of each layer of the point cloud network. The experiment results demonstrate that the optimal fusion depth for pre-trained feature integration occurs at the third layer. At this specific layer, seed points exhibit partial dispersion within the geometric space, enabling the new branch to achieve better multi-modal representations in assimilating rich 2D contextual information into the 3D encoding framework.
IV-E4 Number of Graph Attention Layers
We construct a scene graph and utilize two graph attention layers within the cross-modal interaction module to aggregate node features. The learning objective of the graph network is to capture content related to the referential relationships in textual descriptions by updating node information. However, excessive feature aggregation can result in over-smoothing, which may lead to partial information loss in node representations. Table VI illustrates the effect of varying the number of graph attention layers on overall performance. The results show that better localization outcomes are achieved when the number of layers is two.
IV-E5 CLIP with Different Scales
@@
We construct experiments in Table VII to discuss the effect of different pre-trained CLIP models. All of the models are trained on the large-scale dataset LAION-2B. The network structure of ViT-B/16 is similar to that of ViT-B/32, and the output feature dimensions are both 512, but the image chunks are smaller ($16\times 16$). ViT-L/14 and ViT-H/14 are larger-scale models that output feature vectors of 768 and 1024 dimensions, respectively. The experimental results indicate that scaling up pre-trained models to larger sizes yields no statistically significant improvement in overall performance, suggesting that compact models are sufficient for object-level training requirements. Furthermore, the ViT-B/16 architecture with finer granularity attains a better multi-modal representation compared to the ViT-B/32.
V Conclusion
@@
In this paper, we focus on modeling the referred objects for localization based on referential descriptions in complex scenes and propose a novel framework for 3D visual grounding. The framework first segments scene point clouds into class-agnostic object proposals and then incorporates a pre-trained 2D multi-modal model to enhance 3D visual encoding, enabling the semantic alignment of object features. We construct a semantic-spatial scene graph to facilitate relational learning, where the nodes connect only with semantically related neighbors derived from the textual description. We design an iterative cross-modal interaction module that integrates graph attention layers and vision-language cross-attention layers. Finally, the most relevant object to the description is selected as the target. Experimental results on the Referit3D and ScanRefer benchmarks show that our proposed AS3D delivers better localization performance than other advanced methods, especially demonstrating notable improvement in challenging samples with multiple distractors.

### v3

@@ 4.5.3 Structure of Object Encoder
The object encoder based on the PointNet++ network is a layer-by-layer process of geometric sampling and feature aggregation, so at which layer is it optimal to fuse 2D pre-trained features and generate the new encoding branch? Table 5 illustrates the final localization effect on the set of 2D-enhanced encoding branches across different layers. “layer-ii” denotes that the network architecture bifurcates into two parallel branches at the $i$-th layer, each dedicated to processing distinct feature representations. “att-dim‘’ denotes the mapping dimension of the 2D object features extracted from CLIP before fusion with the point cloud features, to match the output dimension of each layer of the point cloud network. The experiment results demonstrate that the optimal fusion depth for pre-trained feature integration occurs at the third layer. At this specific layer, seed points exhibit partial dispersion within the geometric space, enabling the new branch to achieve better multi-modal representations in assimilating rich 2D contextual information into the 3D encoding framework.
4.5.4 Number of Graph Attention Layers
Table 6: Ablation studies of graph attention layers on Nr3D
@@ 4.5.4 Number of Graph Attention Layers
We construct a scene graph and utilize two graph attention layers within the cross-modal interaction module to aggregate node features. The learning objective of the graph network is to capture content related to the referential relationships in textual descriptions by updating node information. However, excessive feature aggregation can result in over-smoothing, which may lead to partial information loss in node representations. Table 6 illustrates the effect of varying the number of graph attention layers on overall performance. The results show that better localization outcomes are achieved when the number of layers is two.
4.5.5 CLIP with Different Scales
Table 7: Ablation studies of various CLIP models on Nr3D
@@ 4.5.5 CLIP with Different Scales
We construct experiments in Table 7 to discuss the effect of different pre-trained CLIP models. All of the models are trained on the large-scale dataset LAION-2B. The network structure of ViT-B/16 is similar to that of ViT-B/32, and the output feature dimensions are both 512, but the image chunks are smaller ($16\times 16$). ViT-L/14 and ViT-H/14 are larger-scale models that output feature vectors of 768 and 1024 dimensions, respectively. The experimental results indicate that scaling up pre-trained models to larger sizes yields no statistically significant improvement in overall performance, suggesting that compact models are sufficient for object-level training requirements. Furthermore, the ViT-B/16 architecture with finer granularity attains a better multi-modal representation compared to the ViT-B/32.
5 Conclusion
@@
In this paper, we focus on modeling the referred objects for localization based on referential descriptions in complex scenes and propose a novel framework for 3D visual grounding. The framework first segments scene point clouds into class-agnostic object proposals and then incorporates a pre-trained 2D multi-modal model to enhance 3D visual encoding, enabling the semantic alignment of object features. We construct a language-guided scene graph to facilitate relational learning, where the nodes connect only with semantically related neighbors derived from the textual description. We design an iterative cross-modal interaction module that integrates graph attention layers and vision-language cross-attention layers. Finally, the most relevant object to the description is selected as the target. Experimental results on the Referit3D and ScanRefer benchmarks show that our proposed LSVG delivers better localization performance than other advanced methods, especially demonstrating notable improvement in challenging samples with multiple distractors.
Acknowledgments
This work presented in this paper is partially supported by grant from Fundamental Research Funds for the Central Universities(No. 2024ZYGXZR104).